OpenStudy (anonymous):

solve the equation, and check your solution t(4+1)=0

OpenStudy (anonymous):

t(t-4) = 0. You'll have to do like terms here. t*t - 4 = 0 t^2 -4 = 0 Move the -4 to the other side and flip it. t^2 -4+(+4) = 0+(+4) t^2 = 4 Now we do the opposite forms. sqrt(t^2) =4 t = sqrt(4) t = 2 Voila. There's your answer.

OpenStudy (anonymous):

Thanks again you're a lifesaver

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

i think this answer is wrong t(t-4)=0 t*t - t*4 = 0 t^2 - t4 = 0 t^2 = t4 t = 4 that should be the answer if i am not wrong

OpenStudy (anonymous):

Am I seeing wrong or did she change the question? Now I see t(4+1) = 0.

OpenStudy (anonymous):

And she's correct. I forgot all about distributing to the other number. Agh.. That's bad. Don't ever forget to distribute to "everything" that is in the parenthesis. It's a mistake made often.